# Stronger-MAS: Multi-Agent Reinforcement Learning for Collaborative LLMs

### v4

@@ 1 Introduction
Large Language Model (LLM) agents are task-specific workflows (Yao2023ReAct; xi2023rise; Wang2023LLMAgentsSurvey) that utilize LLMs as key components for decision making (shinn2023reflexion), action taking (wang2023voyager), and tool use (Qian2025ToolRL; schick2023toolformer). LLM agents have demonstrated strong promises across various application domains, such as embodied control (ahn2022saycan; wang2023voyager), software engineering (tao2024magis; yu2025orcaloca), expert drug discovery (liu2024drugagent; inoue2024drugagent), and scientific ideation and hypothesis testing (ghafarollahi2024sciagents).
Today, two complementary approaches are widely used to improve the performance of LLM agents: multi-agent systems (MAS) and reinforcement learning (RL). RL treats the LLM as a policy and iteratively updates its weights to strengthen decision-making: at each iteration, the current model interacts with the environment, collects rule-based rewards, and then computes a policy optimization loss to update the parameters (shao2024deepseekmath). In practice, this workflow requires a training stack that supports both scalable rollouts and online updates, e.g., VERL (sheng2025hybridflow) and AReaL (fu2025areal). MAS typically employs prompt-only augmentation on a shared LLM policy for role-based coordination; practical deployments instantiate diverse workflows. Recent studies (belcak2025slm_agentic; chen2024ioa; wang2024moa) further highlight the potential benefits of role-specialized MAS, which adopts distinct models for different roles, enabling role-specialized policies in inference. However, the effectiveness of RL training on role-specialized MAS is underexplored.
@@
A natural next step is to combine the two: using RL to train MAS, such that we gain both stronger learned policies, role-specialized collaboration. However, bringing RL into MAS raises two coupled challenges. First, training a MAS may require concurrently launching multiple models, orchestrating inter-agent environment interactions, and performing independent on-policy parameter updates. But most existing on-policy RL frameworks for LLM agents only support a single model (volcengine2025verl; sheng2024hybridflow; fu2025areal). Second, rollouts from MAS are difficult to group. The advantage must be conditioned on interaction history and role to ensure fair credit assignment. Group-based RL objectives designed for a single agent (volcengine2025verl; Qian2025ToolRL; feng2025group) are not directly applicable to MAS.
@@ 2 Related Work
Role-sharing vs. Role-specialized Policies in MAS. A predominant approach in LLM-based MAS centers on a role-sharing architecture, where a single policy is shared across all agents. In these frameworks, such as AutoGen (wu2023autogen) and MetaGPT (hong2024metagpt), role-specific behavior is elicited at inference time via prompt augmentation. More recently, research has begun to explore role-specialized policies. This shift is motivated by the observation that a single LLM’s performance exhibits significant variance across domains (chen2024ioa; wang2024moa; belcak2025slm_agentic). Consequently, assigning distinct and more suitable models to specialized roles, as demonstrated by ye2025xmas; belcak2025slm_agentic, has emerged as a promising direction for enhancing performance. Despite this architectural evolution, recent surveys (pan2025whydomasfail; guo2024llmma) indicate that most studies focus on inference-time design, leaving the potential of training MAS policies with RL largely underexplored.
RL Training for MAS. RL has become a key technique for LLMs agent training, using group-relative and rule-based rewards to enhance reasoning, long-horizon planning, game, and tool use (feng2025group; wang2025ragen; Qian2025ToolRL; hu2025lmgame). These approaches, however, predominantly operate within a single-agent framework. While a growing body of work attempts to extend RL to Multi-Agent Systems (MAS), most efforts remain confined to limited interaction settings or fixed role structures. For instance, CURE (wang2025cure) focuses on co-evolving a Coder and Unit-Tester using a role-sharing policy specifically for code generation. Similarly, SPIRAL (liu2025spiral) employs self-play in zero-sum games using a single LLM, while MHGPO (chen2025mhgpo) targets retrieval-augmented generation. MAPoRL (park2025maporl; park2025maporl2) and CoRY ma2024cory train LLMs within fixed, homogeneous-role debate workflows. More recent works also exhibit limitations: MARFT Liao2025MARFT  [^1] We compare against MARFT v3, the latest preprint available prior to the completion of this work. restricts agents to single-turn sequential interactions, and MARTI zhang2025marti merely introduces basic single-agent RL algorithms (e.g., GRPO) to the MAS setting. Critically, the implemention for these works is concentrated on the single domain of math. We include a more comprehensive comparison of these related works in Appendix A. In contrast, our study offers a more comprehensive solution. We propose a general algorithm designed for multi-turn, multi-agent environments. Unlike prior works, we conduct a thorough analysis and evaluation of both shared and role-specific policies across diverse MAS workflows and varying domains.
@@ 3 Preliminaries
Following prior work (wang2025cure; ahn2022saycan; Chen2025CodeSteer), we employ domain-specific MAS workflows, as shown in Fig. 2. Our experiments confirm that this prompt-only method outperforms a single-agent baseline (see Tab. 2 and 2 in Sec. 5.2).
@@
Methods for LLM agentic training with group-relative advantages (feng2025group; wang2025ragen; Qian2025ToolRL) operate by first sampling $K$ candidate actions $\{a_{t}^{(c)}\}_{c=1}^{K}$ for a given prompt. Each action is evaluated to obtain a rule-based reward $R(a_{t}^{(c)})$, forming a comparison group: $G=\big\{\,(a_{t}^{(1)},R(a_{t}^{(1)})),\;\ldots,\;(a_{t}^{(K)},R(a_{t}^{(K)}))\,\big\}.$ For each action $a_{t}^{(c)}$ in this group, the relative advantage is then defined as its mean-centered and normalized return.
@@ 4.1 Algorithm Design: AT-GRPO
GRPO’s advantage calculation (Eq. 1) hinges on a fair comparison among all candidates within a group. This fairness is enforced by the reward mechanism itself. As illustrated in Fig. 2 (top), token-level scoring assigns credit to the generated response tokens (Reward Mask=1), while the prompt tokens receive no credit (Reward Mask=0). Since the advantage is determined solely by the quality of the response, a valid and fair comparison is only possible when all responses in a group originate from an identical prompt. Consequently, single-agent LLM-RL methods(wang2025ragen; Qian2025ToolRL; feng2025group) typically form groups by sampling multiple responses to the same question.
In MAS, however, a “prompt” is not only a question description, but also embeds the role-specific context and cross-agent interaction history. For example, in the code tasks depicted in Fig. 2 (a) , where the workflow entails a coder-tester loop: one agent synthesizing code and the other creating unit tests, and they iteratively refine the output until alignment (Fig. 3, top), the turn-2 refinement prompt already contains the turn-1 code, unit tests, and role-specific prompt format, so prompts differ across turns and roles. We therefore adopt agent-wise and turn-wise grouping as a natural extension of tabular-wise group-normalized advantages in GiGPO (feng2025group) to the multi-agent setting: candidates share the same role and turn position, ensuring prompt identity for valid GRPO advantage comparisons.
@@
Inspired by mixed-reward designs in cooperative Multi-Agent RL (Mao2020RewardDesignMARL; Sheikh2020DEMADDPG), we assign credit using a mixture of global and local rewards. At each turn $t$, the environment provides a global team reward $r^{\mathrm{team}}$ and an agent-specific local reward $r^{\mathrm{loc}}_{i}$ that evaluates its subtask performance. These components are combined using a hyperparameter $\alpha$ to form the final reward for agent $i$:
@@ 4.2 MAS Training System
Mainstream RL post-training frameworks for LLMs, e.g., TRL (vonwerra2020trl), VERL (sheng2024hybridflow), AReaL (fu2025areal), and OpenRLHF (hu2024openrlhf) primarily support single-agent RL training, which typically involves: a single agent-environment interaction pattern, a single policy operating on a single data buffer, and a single LLM resource pool. This makes it difficult to (i) train multiple models in on-policy RL, (ii) maintain clean on-policy training data, and (iii) support diverse MAS workflow.
@@
Each policy is managed within an independent resource pool. Following HybridFlow-style  (sheng2025hybridflow), each pool comprises two workers: a RolloutWorker for inference and an UpdateWorker for optimization. During the rollout phase, all policies interact collectively according to the Alg. 1 and MAS workflow; Once collected, each trajectory is routed to the corresponding UpdateWorker, maintaining an on-policy learning regime for every policy.
@@ 5 Experiments
1. Experimental Setup. We train and evaluate Qwen3 models at 1.7B and 8B in the no-thinking mode (Qwen3TR). All runs use a single node with $8\times$ H100 GPUs. The rollout sample size is K=4K{=}4 and the turn horizon is $T=4$ for both multi-agent (MA) and single-agent (SA) settings. The reward-mixing coefficient is $\alpha{=}1$ without further tuning. Full training details appear in Appendix C.1.
@@
We evaluate five variants (all initialized from the same base model): (a) Single Agent (prompt-only): one frozen LLM solves the task end-to-end; (b) Single Agent + GRPO: as (a) but trained with GRPO (shao2024deepseekmath); (c) MAS (prompt-only): role-specialized prompting over a frozen, role-sharing backbone; (d) MAS + RL (role-sharing policy): all roles share one policy and pooled trajectories update it jointly; (e) MAS + RL (role-specialized policies): samples are routed by role and each policy is optimized independently (no parameter sharing).
@@
Sudoku and Sokoban. We evaluate our method on gaming tasks: a $4{\times}4$ Sudoku and a $6{\times}6$ Sokoban. We use instances with an automatic checker, following the symbolic task setup of SymBench (Chen2025CodeSteer). To ensure a fair evaluation, we generate distinct training and validation sets using different random seeds and verify there is no overlap.
Plan-Path. We use a $10{\times}10$ grid-based Plan-Path environment. This follows the checker-backed symbolic task setup in CodeSteer’s SymBench (Chen2025CodeSteer). To separate training and validation, we generate the two splits with distinct random seeds and verify no duplication.
Code Generation. For training, we adopt size-specific corpora: the 1.7B Qwen model is trained on the APPS training split (introductory-difficulty subset) (Hendrycks2021APPS), while the 8B model is trained on CodeContests (DeepMind2024CodeContests). For model-generated code, we use the dataset’s golden unit tests to score correctness; for model-generated UT, we use the dataset’s golden reference solutions to compute the reward. For evaluation, we use three widely adopted coding benchmarks spanning interview-style and contest-style settings: APPS (Hendrycks2021APPS), LiveCodeBench-v6 (white2024livebench), and CodeContests (DeepMind2024CodeContests).
Mathematical Reasoning. We train on the Polaris-Dataset-53K (Polaris2025) and evaluate on several standard mathematical reasoning benchmarks. For validation, we use AIME24/AIME25 (AIME2024; AIME2025) and OlympiadBench (He2024OlympiadBench). All math tasks use verifier-checked numeric scoring.
@@ 5.2 Results and Analysis
MAS + AT-GRPO elevates the success rate from a 14–47% range for the single-agent baseline to 96.0–99.5%. By analyzing the dialogue records between agents, we find this dramatic improvement stems from an emergent collaboration: the tool agent learns to generate correct algorithms (e.g., BFS, $A^{\star}$ search), while the plan agent provides crucial oversight, interpreting execution outcomes and delivering the corrective final action list. On-policy RL training within the MAS enhances inter-agent coordination. Conversely, training agents in isolation results in only limited improvement, as detailed in our ablation study (Sec. 5.4, Tab. 4). Furthermore, on the coding and math benchmarks, our approach yields consistent gains, with absolute gains over the baseline ranging from +2.35 (CodeContests) to +16.30 (APPS) in coding, and from +1.80 (OlympaidBench) to +38.70 (AIME24) in math. We hypothesize two reasons: (1) Base models like Qwen3 have already been extensively trained for these common domains, as noted in their official reports (Qwen3TR), potentially leading to performance saturation. (2) The diverse nature of problems within these domains presents a greater challenge for improvement via RL training.
@@ 5.3 Comparison with other MARL Frameworks
To assess the efficacy of our framework, we conducted ablation studies against representative baselines: As summarized in Tab. 3. we benchmark against three representative baselines: MAPORL (park2025maporl2), MARFT (Liao2025MARFT), and CURE wang2025cure. For fair comparison, we utilize identical base models and dataset splits. Our analysis highlights the advantages of two distinct MAS features: heterogeneous agent roles and multi-turn iterative interaction, as summarized in Tab. 3.
Comparison with MAPORL. We compare our approach with MAPORL using the Phi-3-mini-128k (3.4B) on gsm8k dataset cobbe2021training. While MAPORL relies on a debating mechanism among homogeneous agents, our framework realizes role heterogeneity—synergizing a Reasoning Agent with a Tool-use Agent for verification. This structural specialization proves superior: our untrained MAS achieves 84.4%, outperforming the trained MAPORL (81.0%). With AT-GRPO training, our performance further improves to 88.7%.

### v5

@@ 1 Introduction
Large Language Model (LLM) agents are task-specific workflows (Yao et al., 2023; Xi et al., 2023; Wang et al., 2023b) that utilize LLMs as key components for decision making (Shinn et al., 2023), action taking (Wang et al., 2023a), and tool use (Qian et al., 2025; Schick et al., 2023). LLM agents have demonstrated strong promises across various application domains, such as embodied control (Ahn et al., 2022; Wang et al., 2023a), software engineering (Tao et al., 2024; Yu et al., 2025), expert drug discovery (Liu and others, 2024; Inoue and others, 2024), and scientific ideation and hypothesis testing (Ghafarollahi and Buehler, 2024).
Today, two complementary approaches are widely used to improve the performance of LLM agents: multi-agent systems (MAS) and reinforcement learning (RL). RL treats the LLM as a policy and iteratively updates its weights to strengthen decision-making: at each iteration, the current model interacts with the environment, collects rule-based rewards, and then computes a policy optimization loss to update the parameters (Shao et al., 2024). In practice, this workflow requires a training stack that supports both scalable rollouts and online updates, e.g., VERL (Sheng et al., 2025) and AReaL (Fu and others, 2025). MAS typically employs prompt-only augmentation on a shared LLM policy for role-based coordination; practical deployments instantiate diverse workflows. Recent studies (Belcak et al., 2025; Chen et al., 2024; Wang et al., 2024) further highlight the potential benefits of role-specialized MAS, which adopts distinct models for different roles, enabling role-specialized policies in inference. However, the effectiveness of RL training on role-specialized MAS is underexplored.
@@
A natural next step is to combine the two: using RL to train MAS, such that we gain both stronger learned policies, role-specialized collaboration. However, bringing RL into MAS raises two coupled challenges. First, training a MAS may require concurrently launching multiple models, orchestrating inter-agent environment interactions, and performing independent on-policy parameter updates. But most existing on-policy RL frameworks for LLM agents only support a single model (Volcano Engine, 2025; Sheng et al., 2024; Fu and others, 2025). Second, rollouts from MAS are difficult to group. The advantage must be conditioned on interaction history and role to ensure fair credit assignment. Group-based RL objectives designed for a single agent (Volcano Engine, 2025; Qian et al., 2025; Feng et al., 2025) are not directly applicable to MAS.
@@ 2 Related Work
Role-sharing vs. Role-specialized Policies in MAS. A predominant approach in LLM-based MAS centers on a role-sharing architecture, where a single policy is shared across all agents. In these frameworks, such as AutoGen (Wu et al., 2023) and MetaGPT (Hong et al., 2024), role-specific behavior is elicited at inference time via prompt augmentation. More recently, research has begun to explore role-specialized policies. This shift is motivated by the observation that a single LLM’s performance exhibits significant variance across domains (Chen et al., 2024; Wang et al., 2024; Belcak et al., 2025). Consequently, assigning distinct and more suitable models to specialized roles, as demonstrated by Ye et al. (2025); Belcak et al. (2025), has emerged as a promising direction for enhancing performance. Despite this architectural evolution, recent surveys (Cemri et al., 2025; Guo et al., 2024) indicate that most studies focus on inference-time design, leaving the potential of training MAS policies with RL largely underexplored.
RL Training for MAS. RL has become a key technique for LLMs agent training, using group-relative and rule-based rewards to enhance reasoning, long-horizon planning, game, and tool use (Feng et al., 2025; Wang et al., 2025b; Qian et al., 2025; Hu et al., 2025). These approaches, however, predominantly operate within a single-agent framework. While a growing body of work attempts to extend RL to Multi-Agent Systems (MAS), most efforts remain confined to limited interaction settings or fixed role structures. For instance, CURE (Wang et al., 2025a) focuses on co-evolving a Coder and Unit-Tester using a role-sharing policy specifically for code generation. Similarly, SPIRAL (Liu et al., 2025) employs self-play in zero-sum games using a single LLM, while MHGPO (Chen et al., 2025a) targets retrieval-augmented generation. MAPoRL (Park et al., 2025a; b) and CoRY Ma et al. (2024) train LLMs within fixed, homogeneous-role debate workflows. More recent works also exhibit limitations: MARFT Liao et al. (2025)  [^1] We compare against MARFT v3, the latest preprint available prior to the completion of this work. restricts agents to single-turn sequential interactions, and MARTI Zhang et al. (2025) merely introduces basic single-agent RL algorithms (e.g., GRPO) to the MAS setting. Critically, the implemention for these works is concentrated on the single domain of math. We include a more comprehensive comparison of these related works in Appendix A. In contrast, our study offers a more comprehensive solution. We propose a general algorithm designed for multi-turn, multi-agent environments. Unlike prior works, we conduct a thorough analysis and evaluation of both shared and role-specific policies across diverse MAS workflows and varying domains.
@@ 3 Preliminaries
Following prior work (Wang et al., 2025a; Ahn et al., 2022; Chen et al., 2025c), we employ domain-specific MAS workflows, as shown in Fig. 2. Our experiments confirm that this prompt-only method outperforms a single-agent baseline (see Tab. 2 and 2 in Sec. 5.2).
@@
Methods for LLM agentic training with group-relative advantages (Feng et al., 2025; Wang et al., 2025b; Qian et al., 2025) operate by first sampling $K$ candidate actions $\{a_{t}^{(c)}\}_{c=1}^{K}$ for a given prompt. Each action is evaluated to obtain a rule-based reward $R(a_{t}^{(c)})$, forming a comparison group: $G=\big\{\,(a_{t}^{(1)},R(a_{t}^{(1)})),\;\ldots,\;(a_{t}^{(K)},R(a_{t}^{(K)}))\,\big\}.$ For each action $a_{t}^{(c)}$ in this group, the relative advantage is then defined as its mean-centered and normalized return.
@@ 4.1 Algorithm Design: AT-GRPO
GRPO’s advantage calculation (Eq. 1) hinges on a fair comparison among all candidates within a group. This fairness is enforced by the reward mechanism itself. As illustrated in Fig. 2 (top), token-level scoring assigns credit to the generated response tokens (Reward Mask=1), while the prompt tokens receive no credit (Reward Mask=0). Since the advantage is determined solely by the quality of the response, a valid and fair comparison is only possible when all responses in a group originate from an identical prompt. Consequently, single-agent LLM-RL methods(Wang et al., 2025b; Qian et al., 2025; Feng et al., 2025) typically form groups by sampling multiple responses to the same question.
In MAS, however, a “prompt” is not only a question description, but also embeds the role-specific context and cross-agent interaction history. For example, in the code tasks depicted in Fig. 2 (a) , where the workflow entails a coder-tester loop: one agent synthesizing code and the other creating unit tests, and they iteratively refine the output until alignment (Fig. 3, top), the turn-2 refinement prompt already contains the turn-1 code, unit tests, and role-specific prompt format, so prompts differ across turns and roles. We therefore adopt agent-wise and turn-wise grouping as a natural extension of tabular-wise group-normalized advantages in GiGPO (Feng et al., 2025) to the multi-agent setting: candidates share the same role and turn position, ensuring prompt identity for valid GRPO advantage comparisons.
@@
Inspired by mixed-reward designs in cooperative Multi-Agent RL (Mao et al., 2020; Sheikh and Bölöni, 2020), we assign credit using a mixture of global and local rewards. At each turn $t$, the environment provides a global team reward $r^{\mathrm{team}}$ and an agent-specific local reward $r^{\mathrm{loc}}_{i}$ that evaluates its subtask performance. These components are combined using a hyperparameter $\alpha$ to form the final reward for agent $i$:
@@ 4.2 MAS Training System
Mainstream RL post-training frameworks for LLMs, e.g., TRL (von Werra et al., 2020), VERL (Sheng et al., 2024), AReaL (Fu and others, 2025), and OpenRLHF (Hu et al., 2024) primarily support single-agent RL training, which typically involves: a single agent-environment interaction pattern, a single policy operating on a single data buffer, and a single LLM resource pool. This makes it difficult to (i) train multiple models in on-policy RL, (ii) maintain clean on-policy training data, and (iii) support diverse MAS workflow.
@@
Each policy is managed within an independent resource pool. Following HybridFlow-style  (Sheng et al., 2025), each pool comprises two workers: a RolloutWorker for inference and an UpdateWorker for optimization. During the rollout phase, all policies interact collectively according to the Alg. 1 and MAS workflow; Once collected, each trajectory is routed to the corresponding UpdateWorker, maintaining an on-policy learning regime for every policy.
@@ 5 Experiments
1. Experimental Setup. We train and evaluate Qwen3 models at 1.7B and 8B in the no-thinking mode (Yang and the Qwen Team, 2025). All runs use a single node with $8\times$ H100 GPUs. The rollout sample size is K=4K{=}4 and the turn horizon is $T=4$ for both multi-agent (MA) and single-agent (SA) settings. The reward-mixing coefficient is $\alpha{=}1$ without further tuning. Full training details appear in Appendix C.1.
@@
We evaluate five variants (all initialized from the same base model): (a) Single Agent (prompt-only): one frozen LLM solves the task end-to-end; (b) Single Agent + GRPO: as (a) but trained with GRPO (Shao et al., 2024); (c) MAS (prompt-only): role-specialized prompting over a frozen, role-sharing backbone; (d) MAS + RL (role-sharing policy): all roles share one policy and pooled trajectories update it jointly; (e) MAS + RL (role-specialized policies): samples are routed by role and each policy is optimized independently (no parameter sharing).
@@
Sudoku and Sokoban. We evaluate our method on gaming tasks: a $4{\times}4$ Sudoku and a $6{\times}6$ Sokoban. We use instances with an automatic checker, following the symbolic task setup of SymBench (Chen et al., 2025c). To ensure a fair evaluation, we generate distinct training and validation sets using different random seeds and verify there is no overlap.
Plan-Path. We use a $10{\times}10$ grid-based Plan-Path environment. This follows the checker-backed symbolic task setup in CodeSteer’s SymBench (Chen et al., 2025c). To separate training and validation, we generate the two splits with distinct random seeds and verify no duplication.
Code Generation. For training, we adopt size-specific corpora: the 1.7B Qwen model is trained on the APPS training split (introductory-difficulty subset) (Hendrycks et al., 2021), while the 8B model is trained on CodeContests (DeepMind, 2024). For model-generated code, we use the dataset’s golden unit tests to score correctness; for model-generated UT, we use the dataset’s golden reference solutions to compute the reward. For evaluation, we use three widely adopted coding benchmarks spanning interview-style and contest-style settings: APPS (Hendrycks et al., 2021), LiveCodeBench-v6 (White et al., 2024), and CodeContests (DeepMind, 2024).
Mathematical Reasoning. We train on the Polaris-Dataset-53K (An et al., 2025) and evaluate on several standard mathematical reasoning benchmarks. For validation, we use AIME24/AIME25 (Mathematical Association of America & AoPS Community, 2024; 2025) and OlympiadBench (He et al., 2024). All math tasks use verifier-checked numeric scoring.
@@ 5.2 Results and Analysis
MAS + AT-GRPO elevates the success rate from a 14–47% range for the single-agent baseline to 96.0–99.5%. By analyzing the dialogue records between agents, we find this dramatic improvement stems from an emergent collaboration: the tool agent learns to generate correct algorithms (e.g., BFS, $A^{\star}$ search), while the plan agent provides crucial oversight, interpreting execution outcomes and delivering the corrective final action list. On-policy RL training within the MAS enhances inter-agent coordination. Conversely, training agents in isolation results in only limited improvement, as detailed in our ablation study (Sec. 5.4, Tab. 4). Furthermore, on the coding and math benchmarks, our approach yields consistent gains, with absolute gains over the baseline ranging from +2.35 (CodeContests) to +16.30 (APPS) in coding, and from +1.80 (OlympaidBench) to +38.70 (AIME24) in math. We hypothesize two reasons: (1) Base models like Qwen3 have already been extensively trained for these common domains, as noted in their official reports (Yang and the Qwen Team, 2025), potentially leading to performance saturation. (2) The diverse nature of problems within these domains presents a greater challenge for improvement via RL training.
@@ 5.3 Comparison with other MARL Frameworks
To assess the efficacy of our framework, we conducted ablation studies against representative baselines: As summarized in Tab. 3. we benchmark against three representative baselines: MAPORL (Park et al., 2025b), MARFT (Liao et al., 2025), and CURE Wang et al. (2025a). For fair comparison, we utilize identical base models and dataset splits. Our analysis highlights the advantages of two distinct MAS features: heterogeneous agent roles and multi-turn iterative interaction, as summarized in Tab. 3.
Comparison with MAPORL. We compare our approach with MAPORL using the Phi-3-mini-128k (3.4B) on gsm8k dataset Cobbe et al. (2021). While MAPORL relies on a debating mechanism among homogeneous agents, our framework realizes role heterogeneity—synergizing a Reasoning Agent with a Tool-use Agent for verification. This structural specialization proves superior: our untrained MAS achieves 84.4%, outperforming the trained MAPORL (81.0%). With AT-GRPO training, our performance further improves to 88.7%.